1.1. Elkins, a manufacturer of ice makers, realizes a cost of $300 for every unit it produces. Its total fixed costs equal S10 million. If the company manufactures 500,000 units compute the following: a. unit cost markup price if the company desires a 15% return on sales c. Rol price if the company desires a 20% return on an investment of $1 million 1.2. A consumer purchases a coffee maker from a retailer for S150. The retailers markup is 25%, and the wholesalers markup is 30%, both based on selling price. For what price does the manufacturer sell the product to the wholesaler? 1.3. A lawnmower manufacturer has a unit cost of S150 and wishes to achieve a margin of 20% based on selling price. If the manufacturer sells directly to a retailer who then adds a set margin of 25% based on selling price, determine the retail price charged to consumers.
Answer #1

1.1

Given That,

Variable Cost Per unit = 300

Total Fixed Cost = 10000000

Total Unit Manufactured = 50000 units

a.

Unit Cost = Variable Cost Per Unit + (Total Fixed Cost/Total Unit Manufactured)

Unit Cost = 300 + (10000000/50000)

Unit Cost = $500 per unit

b.

Markup Price = Unit Cost/(1-Desired Return on Sales)

Markup Price = 500/(1-0.15) = 500/0.85 = 588.24

c.

Markup Price = Unit Cost + (Return on investment*Investment/Unit Sales)

Markup Price = 500 + (0.20*1000000/50000) = 500 +4 = 504